Three numbers are in G.P their sum is 7 and sum of their squares in 21 find the numbers

Grade:12

2 Answers

Shona
8 years ago
b^2=ac-----(1)
a+b+c=7----(2)
a^2+b^2+c^2=21------(3)
solve the three equations
you will get
a=1    &     a=4
b=2    &     b=2
c=4    &     c=1
Gavvala Ganesh
8 years ago
we know it was in g.p so,b^22=ac------eq(1)
Given that
a+b+c=7----(2)
a^2+b^2+c^2=21=---(3)
solve we get
a=1 and a=4
b = 2 and b = 2
c = 4 and c = 1.